$10,000 in two mutual funds, one rose 6% in one year and one rose 9% in one year. It rose a total of $684 in one year. How much was invested in each mutual fund?
Answer by rmromero(383)

$10,000 in two mutual funds, one rose 6% in one year and one rose 9% in one year. It rose a total of $684 in one year. How much was invested in each mutual fund?
What is ask?
How much was invested in each mutual fund?


Given:
the two mutual fund is $ 10, 000
One fund rose 6% in one year
The other rose 9% in one year
The total rose of the funds is $ 684


Representation:
Let x = fund with 6% interest
y = fund with 9% interest


Equation:
x + y = $ 10, 000
0.06x + 0.09y = $ 684


Solve:
Multiply 100 to 0.06x + 0.09y = $ 684 to eliminate decimals
x + y = 10, 000
6x + 9y = 68, 400


Multiply -6 to x + y = 10, 000 to eliminate x
-6x - 6y = -60, 000
6x + 9y = 68, 400


Add:
3y = 8, 400
y = $ 2, 800


Solve for x:
x + y = 10, 000
x + 2, 800 = 10, 000
x = $ 7, 200
Checking:
x + y = 10,000
7, 200 + 2, 800 = 10,000
10,000= 10,000 ----------->> true
0.06x + 0.09y = 684
432 + 252 = 584
684 = 684 ----------->> true


Therefore the fund where $2,800 for 9% interest and $7,200 for 6% interest